OpenStudy (anonymous):

Please Help! What are the x-intercepts of y = 2(x - 4)(x + 1)

OpenStudy (anonymous):

let y=0 ans solve for x

OpenStudy (anonymous):

(0,-4) that should be your answer

OpenStudy (whpalmer4):

@weasle the function in question is a parabola which crosses the x-axis twice. You're trying to find the values of \(x\) where that happens.

OpenStudy (anonymous):

@weasle its not one of the choices

OpenStudy (anonymous):

hold on

OpenStudy (whpalmer4):

\[0 = 2(x-4)(x+1)\]Your job is to find the values of \(x\) that make that statement true. You should be able to do this just by thinking a little bit.

OpenStudy (whpalmer4):

If you have \[a*b = 0\]what must be true about a and b?

OpenStudy (whpalmer4):

One of them has to be 0, right?

OpenStudy (anonymous):

its -1 and 4 thank you all for trying to help

OpenStudy (whpalmer4):

Don't go, you're not done yet.

OpenStudy (anonymous):

@whpalmer4 ok what else needs to be done?

OpenStudy (whpalmer4):

You need to understand how to get the answer :-) So the only way we can get \[0 = 2(x-4)(x+1)\] is if \[x-4=0\]or \[x+1=0\]If you solve those two equations, you'll get the answer values. Usually, you'll be given an equation that isn't already factored, and then finding the solutions is not as easy. For example, it's obvious that you can look at this one and say "x = 4, x = -1" but if they gave us the equation multiplied out, wouldn't be so easy to see that for \[2x^2-6x-8=0\]you get the same answers, would it?

OpenStudy (whpalmer4):

But you do: \[2(4)^2-6(4)-8 = 0\]\[2*16-24-8 = 0\]\[32-24-8=0\]\[0=0\] and \[2(-1)^2-6(-1)-8=0\]\[2(1)+6-8 = 0\]\[2+6-8=0\]\[0=0\]

OpenStudy (whpalmer4):

You'll be seeing a lot of this concept, trust me :-)

OpenStudy (anonymous):

@whpalmer4 Thank you so much for teaching me how to check my answers :)

OpenStudy (anonymous):

sorry for not being able to help more. :(

OpenStudy (anonymous):

@weasle Its ok thank you for trying :)

OpenStudy (whpalmer4):

I figure if the problem is worth doing, you might as well make sure you've got the right answer :-)

OpenStudy (anonymous):

@whpalmer4 Thank you!

OpenStudy (whpalmer4):

Here's a graph of that function, by the way. You can clearly see the solutions are the points where the parabola crosses the x axis.